# Know your exceptions

### 01 Mar 2024 v1

@@ 3.3 Characterising Exceptions
In the previous sections we set the ground for comparing the different non-monotonic logics in order to have insights on the commitment they do which could be relevant for a modeller who would like to use them. The key point emerged in the latter subsection: given the relationship between universal and defeasible generalisations and since the aim of non-monotonic logics is to represent defeasible generalisations, we can compare them according to how they try to characterise the exceptions.
According to this framework we can propose a new definition of exception:
Definition 3
@@
Firstly, individuating exceptionality as a relation between an individual and a generalisation is obviously preserved. The fact that it is excluded from that generalisation is also preserved since the idea is exactly that the subset to which the exception belong is considered as the set of completers, which if mentioned in the generalisation it allows to individuate the exact subset in the scope of the defeasible generalisation. It also captures the idea that exceptions do not falsify the generalisation, since they do not falsify the defeasible generalisation, but they still have the property contradicting the generalisation since they excluded from the subset instantiating the generalised property.
However, the definition above leaves unspecified the idea of justification for the exclusion of the exceptions and of the different kinds of knowledge as their sources: these are exactly those factors that allow to individuate actual exceptions, which is the goal of the specific formalisations. Therefore, we can see this definition as incomplete: studying how each non-monotonic logics specifies this idea of justification is exactly our interest.
In the next section we will explain four formalisms in their key aspects, in order to see how they can complete the definition of exception above and so what does it mean in those cases to be an exception and therefore how defeasibility is understood.
@@ 4 Formal Models of Defeasible Reasoning
Here we give a brief sketch of the approaches in order to make the comparison in the next section understandable, for a more precise and formal introduction we refer to [2]. We provide a brief introduction to the approaches in Appendix A.
@@ 4.2 Circumscription
The key idea of circumscription is introducing a class abnormality predicates $Ab_{i}$, which is minimised. Therefore, we can represent $\widetilde{\forall}$-generalisation as $\forall x((P(x)\land\lnot Ab(x))\rightarrow Qx)$ which can be interpreted as if something is a P and is not abnormal then it is a Q too.
@@
For example, consider the knowledge base containing the propositions $\forall x(Bird(x)\land\lnot Ab(x)\rightarrow Flies(x))$, $Bird(tweety)$, $Bird(chilly)$, $\lnot Flies(chilly)$, $(tweety\neq chilly)$. Chilly is correctly considered abnormal, since it does not fly and we can conclude that Tweety flies because we consider the minimal models, and a model where $Ab$ contains only Chilly is more normal than one containing also Tweety.
@@ 4.3 Default Logic
For instance, consider the knowledge base containing the propositions $Bird(tweety)$, $Bird(chilly)$, $\lnot Flies(chilly)$ and the rule $Birds(x):Flies(x)/Flies(x)$. In this case we have only a rule which says that if something is a bird and it is not inconsistent to assume that it flies, then we can conclude that it flies. We can see that this is true only for Tweety, since we know that Chilly does not fly and so we have in our expansion that Tweety flies.
@@ 4.4 Autoepistemic Logic
We introduce the unary operator B with the meaning of “it is believed that…”.
For instance, imagine our knowledge base containing the following sentences $Bird(tweety)$, $Bird(chilly)$, $(tweety\neq chilly)$, $\lnot Flies(chilly)$, $\forall x(Bird(x)\land\lnot\textbf{B}\lnot Flies(x)\rightarrow Flies(x))$. From this KB, since we have $\lnot Flies(chilly)$ and $\lnot\textbf{B}\lnot Flies(chilly)$, we have a $\textbf{B}\lnot Flies(chilly)$ because there is no way of entailing $\lnot Flies(x)$. Therefore, we cannot infer $Flies(chilly)$ but we can infer $Flies(tweety)$.
Now that we have an idea of the key point of these four theories we can start to compare them and show in which regard they differ.
@@ 5 The Comparison
The comparison of the four formalisms is developed along four lines: if the defeasibility is represented through a syntactic or a semantic mechanism; if defeasibility is understood at the epistemic or at the ontological level; if exceptions are represented explicitly or implicitly; and finally if the defeasible generalisations are understood as formulae representing propositions or as meta-language rules.
In table 1 we can see an overview of the results of the comparison, which we discuss in the next subsections.
Semantic Vs Syntactic Epistemic Vs Ontological Explicit Vs Implicit Logical vs Meta-Logical
@@ 5.1 Syntactic or Semantic Approach
As we have said above, having CWA in this comparison is not completely fair, since we do not really have a representation of defeasibility in the system. However, it may be interesting for some discussion that we can have regarding this system. In fact, even if the CWA is a semantic assumption, because teh assumption regards the world, the mechanism used is syntactical, since what counts is the presence or not of literals in the KB.
Regarding circumscription, the situation is more simple since we introduce a new predicate with an interpretation that uses an order on models. This makes it a semantic approach for defeasibility. Default and autoepistemic logic, instead, are both syntactical systems. In fact, the default rules do not have an interpretation but they work on the basis of the notion of syntactical consistency. Whereas, the introspection operator has not a semantic interpretation and so all the non-monotonic mechanism rely on syntactic basics.
@@ 5.2 Epistemic or Ontological Level
The second important distinction regards the source of the defeasibility, that is if the system represent defeasibility as something due to how the world is or to how we represent it. In the former case, we are considering defeasibility as something occurring at the ontological level, while in the latter we are considering it as something at the level of knowledge, that is epistemic.
This distinction seems related with the previous one, in fact, in this comparison they overlap. However, since the distinctions are in principle different, since the former regards the type of mechanism used for modelling defeasibility, whereas the latter consider at which level defeasibility emerges and is understood, we maintain them separate and we will investigate the correlation as future work.
In CWA, despite its name, defeasibility comes from our way of knowing, since the assumption is that we know all the relevant knowledge. Therefore, here we are at the epistemic level.
In circumscription, we are at the ontological level instead. In fact, defeasibility comes from the fact that we recognise abnormal individuals, to which generalisations regarding normal circumstances do not apply.
In default logic, the interpretation of the KB, composed of the set of FOL sentences and the set of default rules, is that of a theory and the key notion is that of consistency among the beliefs entailed by the theory. Therefore ,here we are treating defeasibility as something occurring at the epistemic theory.
In the case of autoepistemic logic it is not immediate the level at which we are considering the defeasibility. In fact, at the object language we are working with beliefs, but the new introspection operator is not a predicate, therefore it is not at the object level. So, since the defeasible mechanism is realised through the definition of extensions like for default logic, the predominant level is that of epistemology.
@@ 5.3 Explicit or Implicit Representation of Exceptions
The third line of comparison which allows to distinguish the approaches regards how exceptions are represents in the system, more precisely if the are represented explicitly or implicitly. This is relevant, since having in the ontology the class, or classes, of exceptions would make a big difference with respect of not having the. In fact, on one hand, having explicit exceptions in the model would mean that exceptional entities would be immediately recognisable and can be treated also with respect of their being exceptions. On the other hand, this would also mean that the ontology should be structured in order to accommodate also the property of being an exception.
In this distinction and the next one, it emerges the peculiarity of CWA. In fact, since there is not defeasible generalisations, in CWA there is not a notion of exception and a fortiori there is not a notion of implicit of explicit exception.
Circumscription, instead, is probably the easiest case for this distinction, in fact, here are considered exceptions those individuals which necessarily are abnormal. This is due to the mechanism of reducing the extension of $Ab$ as much as possible. Therefore, in this system exceptions are represented explicitly as those individuals which are in one of the abnormal predicates.
In default logic, exceptions are individuate exploiting inconsistency. That is, the instances which are exceptions are those which satisfy the prerequisite, but not the justification and therefore, assuming them would lead to an inconsistency. However, these instances are not marked in some way, so to recognise them we have to look at the default rule and search is there are individuals which satisfy the prerequisite but not the justification. Consequently, exceptions here are represented implicitly.
The case of autoepistemic logic is similar to that of default logic: we can consider exceptions those beliefs about individuals which do not correspond to the introspection status expressed through the operator in the antecedent. Which means that they satisfy the classical part of the antecedent but not the one with the introspective operator, which can be seen respectively as the analogous of the prerequisite and the justification of a default rule. Therefore, autoepistemic logic represent exceptions implicitly too.
@@ 5.4 Logical or Meta-logical
The last comparison regards the form given to defeasible generalisations, that is if they are modelled at the logical level and so we can reason on them as we do with classical FOL formulas or if they are outside the language and so they are meta-logical. This distinction is interesting because it allows to distinguish the systems where it is possible to exploit the representations of the defeasible generalisations to reason on them and those where it is not possible to reason about them.
Again, this distinction cannot be applied to CWA since it does not represent defeasible generalisations. For what regards circumscription instead, defeasible generalisations are modelled at the logical level, whereas in default logic, as suggested above, they are considered rules at the meta-logical.
More interesting is the case of autoepistemic logic. In this case, defeasible generalisations are formulae in the language, therefore they are treated at the logical level. However, they do not have a standard extensional semantic, since the introspection operator gain its meaning through the notion of stable expansion, which is a syntactic notion. Therefore, even if autoepistemic logic can be classified as modelling defeasible generalisations as formulae on which we can reason about, it does so in a peculiar way with respect of circumscription.
@@ 6 Conclusions and Future Works
In this paper we introduced a framework based on the notions of generalisation and exception for a comparison of non-monotonic logics that should help a modeller in choosing the better approach according to her needs for the specific domain. In section 2 we analysed and discuss what does it mean to be exceptional from a broader point of view. This discussion supplied the basis for a more precise discussion in section 3 thanks also to the insights coming from the study of generics and ct-laws. In this section we developed the definitions for the key notions of defeasible generalisation and exception for the framework. Then, after having described the formal approaches used in the comparison in section 4, we did the comparison in section 5, which is summarised in table 1.
As future works, we plan to develop further the framework for example through a comparison with [19], where an idea of generalised quantifier is discussed for representing defeasible reasoning.
Moreover, we want to extend the comparison to other kinds of non-monotonic logics like preferential approaches [20, 21] and modal approaches [16], but also to application of non-monotonic logics in the context of description logics like in [8, 22].
Another interesting discussion we would like to develop is the possibility of representing defeasible reasoning through monotonic logics.
Finally, we plan to exploit the analysis developed her on the notion of exceptions and generalisations to refine our formal system outlined in [10].
@@ Appendix A Formal Models of Defeasible Reasoning: extended
In this appendix we present in an extended version the four approaches discussed above: Closed World Assumption (henceforth CWA), Circumscription, Default Logic and Autoepistemic Logic.

### 05 Mar 2024 v2

@@ 3.3 Characterising Exceptions
In the previous sections we set the ground for comparing the different non-monotonic logics in order to have insights on the commitments they make which could be relevant for a modeller who would like to use them. The key point emerged in the latter subsection: given the relationship between universal and defeasible generalisations and since the aim of non-monotonic logics is to represent defeasible generalisations, we can compare them according to how they try to characterise the exceptions.
According to this framework we can propose a new understanding of exception:
Concept Explication. 3
@@
Firstly, individuating exceptionality as a relation between an individual and a generalisation is obviously preserved. The fact that it is excluded from that generalisation is also preserved since the idea is exactly that the subset to which the exception belongs is considered as the set of completers, which if mentioned in the generalisation allows to individuate the exact subset in the scope of the defeasible generalisation. It also captures the idea that exceptions do not falsify the generalisation, since they do not falsify the defeasible generalisation, but they still have the property contradicting the generalisation since they are excluded from the subset instantiating the generalised property.
However, the definition above leaves unspecified the idea of justification for the exclusion of the exceptions and of the different kinds of knowledge as their sources: these are exactly those factors that allow to individuate actual exceptions, which is the goal of the specific formalisations. Therefore, we can see this definition as incomplete: studying how each non-monotonic logic specifies this idea of justification is exactly our interest.
In the next section, we will explain four formalisms considering their key aspects in order to see how they can complete the definition of exception above, and thus to understand what it means in each case to be an exception, and therefore how defeasibility is understood.
@@ 4 Formal Models of Defeasible Reasoning
Here we give a sketch of the approaches in order to make the comparison in the next section understandable. We provide a brief introduction to the approaches in Appendix A and for a more precise and formal discussion we refer to [2].
@@ 4.1 Closed World Assumption
The central idea of the CWA is that what is not explicitly mentioned can reasonably be assumed to be false. So this leads to defeasibility because you could come to know something new, which changes it from being considered false to being considered true. However, CWA is a defeasible system in a broad sense. In fact, it does not represent defeasible knowledge at the object level, rather the modeller can update manually the KB and so change information in it. So to say, with the CWA we are always in the default state of knowledge and we cannot really represent non-default cases.
@@ 4.2 Circumscription
The key idea of circumscription is introducing a class of abnormality predicates $Ab_{i}$, which is minimised. Therefore, we can represent $\widetilde{\forall}$-generalisation as $\forall x((P(x)\land\lnot Ab(x))\rightarrow Qx)$ which can be interpreted as “if $x$ is a $P$ and is not abnormal, then it is a $Q$ too”.
@@
For example, consider the knowledge base containing the propositions: $\forall x(Bird(x)\land\lnot Ab(x)\rightarrow Flies(x))$, $Bird(tweety)$, $Bird(chilly)$, $\lnot Flies(chilly)$, $(tweety\neq chilly)$. Chilly is correctly considered abnormal, since it does not fly and we can conclude that Tweety flies because we consider the minimal models, and a model where $Ab$ contains only Chilly is more normal than one containing also Tweety.
@@ 4.3 Default Logic
The intuition behind default logic is to add a set of rules that allow to specify which assumptions should be made, given that they are consistent. A rule has the following form: $\langle\alpha:\beta/\delta\rangle$ where $\alpha$ is a sentence corresponding to the prerequisite for the application of the rule, $\beta$ is one that corresponds to the justification for the application of the rule and $\gamma$ is the conclusion of the rule. The idea is that if we know $\alpha$ and $\beta$ is consistent with what we know, we can conclude $\delta$.
These rules are the ones that allow to represent defeasible generalisations. In fact, default rules do not necessarily apply to all individuals, but only to those which are consistent to do so. Therefore, exceptions to a default rule are those individuals for which the justification does not apply. Interestingly, this is an inverse perspective with respect to our view above, since in this case the justification is needed for applying the defeasible generalisation, rather than to individuate the exceptions.
For instance, consider the knowledge base containing the propositions: $Bird(tweety)$, $Bird(chilly)$, $\lnot Flies(chilly)$ and the rule $Birds(x):Flies(x)/Flies(x)$. In this case we have only a rule which says that if something is a bird and it is not inconsistent to assume that it flies, then we can conclude that it flies. We can see that this is true only for Tweety, since we know that Chilly does not fly and so we have in our expansion that Tweety flies.
@@ 4.4 Autoepistemic Logic
In autoepistemic logic, the intuition is that of trying to exploit introspection, represented with the introduction of an unary connective which is used to represent what is believed. This allows to represent defeasibility because adding or removing beliefs affect the beliefs I believe, in this way the conclusions that I reach by exploiting what I know to believe can change. That is, we can exploit the indexicality of the introspection operator [19].
@@
We introduce the unary operator B with the meaning of “it is believed that…”. For instance, imagine our knowledge base containing the following sentences $Bird(tweety)$, $Bird(chilly)$, $(tweety\neq chilly)$, $\lnot Flies(chilly)$, $\forall x(Bird(x)\land\lnot\textbf{B}\lnot Flies(x)\rightarrow Flies(x))$. From this KB, since we have $\lnot Flies(chilly)$ and $\lnot\textbf{B}\lnot Flies(chilly)$, we have a $\textbf{B}\lnot Flies(chilly)$ because there is no way of entailing $\lnot Flies(x)$. Therefore, we cannot infer $Flies(chilly)$ but we can infer $Flies(tweety)$.
@@ 5 The Comparison
Now that we have an idea of the key points of these four theories we can start to compare them and show in which regard they differ. The comparison of the four formalisms is developed along four lines: if the defeasibility is represented through a syntactic or a semantic mechanism; if defeasibility is understood at the epistemic or at the ontological level; if exceptions are represented explicitly or implicitly; and finally, if the defeasible generalisations are understood as formulae representing propositions (at the logical level) or as meta-language rules (at the meta-logical level). In Table 1, we show an overview of the results of this comparison, which we will discuss in the next subsections.
Semantic Vs Syntactic
Epistemic Vs Ontological
Explicit Vs Implicit
Logical vs Meta-Logical
@@ 5.1 Syntactic or Semantic Approach
As we have said above, having CWA in this comparison is not completely fair, since we do not really have a representation of defeasibility in the system. However, it may be interesting for some discussion that we can have regarding this system. In fact, even if the CWA is a semantic assumption, because the assumption concerns the world, the mechanism used is syntactical, since what counts is the presence or not of literals in the KB.
Regarding circumscription, the situation is more simple since we introduce a new predicate with an interpretation that uses an order on models. This makes it a semantic approach for defeasibility. Default and autoepistemic logic, instead, are both syntactical systems. In fact, the default rules do not have an interpretation but they work on the basis of the notion of syntactical consistency. Whereas, the introspection operator has not a semantic interpretation and so the entire non-monotonic mechanism relies on a syntactic basis.
@@ 5.2 Epistemic or Ontological Level
The second important distinction concerns the source of the defeasibility, that is if the system represents defeasibility as something due to how the world is or to how we represent it. In the former case, we are considering defeasibility as something occurring at the ontological level, while in the latter we are considering it as something at the level of knowledge, that is epistemic. This distinction seems related to the previous one, in fact, in this comparison they overlap. However, since the distinctions are in principle different, since the former concerns the type of mechanism used for modelling defeasibility, whereas the latter considers at which level defeasibility emerges and is understood, we maintain them separately and will investigate their correlation as future work. In CWA, despite its name, defeasibility comes from our way of knowing, since the assumption is that we know all the relevant knowledge. Therefore, here we are at the epistemic level. In circumscription, we are at the ontological level. In fact, defeasibility comes from the fact that we recognise abnormal individuals, to which generalisations regarding normal circumstances do not apply. In default logic, the interpretation of the KB, composed of the set of FOL sentences and the set of default rules, is that of a theory and the key notion is that of consistency among the beliefs entailed by the theory. Therefore, here we are treating defeasibility as something occurring at the epistemic theory. In the case of autoepistemic logic, it is not immediate the level at which we are considering the defeasibility. In fact, at the object language we are working with beliefs, but the new introspection operator is not a predicate, therefore it is not at the object level. So, since the defeasible mechanism is realised through the definition of extensions like for default logic, the predominant level is that of epistemology.
@@ 5.3 Explicit or Implicit Representation of Exceptions
The third line of comparison which allows to distinguish the approaches regards how exceptions are represented in the system, more precisely if they are represented explicitly or implicitly. Having in the ontology a class (or classes) of exceptions is a relevant issue. In fact, on one hand, having explicit exceptions in the model would mean that exceptional entities would be immediately recognisable and can be treated also with respect of their being exceptions. On the other hand, this would also mean that the ontology should be structured in order to accommodate the property of being an exception.
In this distinction and the next one, it emerges the peculiarity of CWA. In fact, since there are no defeasible generalisations, in CWA there is not a notion of exception and a fortiori there is not a notion of implicit of explicit exception.
Circumscription, instead, is probably the easiest case for this distinction: in fact, in this formalism exceptions are those individuals which necessarily are abnormal. This is due to the mechanism of reducing the extension of $Ab$ as much as possible. Therefore, in this system exceptions are represented explicitly as those individuals which are in one of the abnormal predicates.
In default logic, exceptions are identified by inconsistency. That is, the instances which are exceptions are those which satisfy the prerequisite, but not the justification and therefore, assuming them would lead to an inconsistency. However, these instances are not marked in some way, so to recognise them we have to consider the default rules and look for those individuals which satisfy the prerequisite but not the justification. Consequently, exceptions here are represented implicitly.
The case of autoepistemic logic is similar to that of default logic: we can consider exceptions those beliefs about individuals which do not correspond to the introspection status expressed through the operator in the antecedent. This means that they satisfy the classical part of the antecedent but not the one with the introspective operator, which can be seen respectively as the analogous of the prerequisite and the justification of a default rule. Therefore, autoepistemic logic represents exceptions implicitly too.
@@ 5.4 Logical or Meta-logical
The last comparison regards the form given to defeasible generalisations, that is if they are modelled at the logical level and so we can reason on them as we do with classical FOL formulas or if they are outside the language and so they are meta-logical. This distinction is interesting because it allows us to distinguish the systems where it is possible to exploit the representations of the defeasible generalisations to reason on them and those where it is not possible to reason about them.
Again, this distinction cannot be applied to CWA since it does not represent defeasible generalisations. For what regards circumscription instead, defeasible generalisations are modelled at the logical level, whereas in default logic, as suggested above, they are considered rules at the meta-logical level.
More interesting is the case of autoepistemic logic. In this case, defeasible generalisations are formulae in the language, therefore they are treated at the logical level. However, they do not have a standard extensional semantics since the introspection operator gains its meaning through the notion of stable expansion, which is a syntactic notion. Therefore, even if autoepistemic logic can be classified as modelling defeasible generalisations as formulae about which we can reason, it does so in a peculiar way with respect of circumscription.
@@ 6 Conclusions and Future Works
In this paper, we introduced a framework based on the notions of generalisation and exception for a comparison of non-monotonic logics that should help a modeller in choosing the better approach according to her needs for a specific domain. In Section 2 we analysed and discussed what it means to be exceptional from a broader point of view. This discussion supplied the basis for a more precise discussion in Section 3 thanks also to the insights coming from the study of generics and ct-laws. In this section we developed the definitions for the key notions of defeasible generalisation and exception for the framework. Then, after having described the formal approaches used in the comparison in Section 4, we did the comparison in Section 5, which is summarised in Table 1.
As future works, we plan to develop further the framework for example through a comparison with [20], where an idea of generalised quantifier is discussed for representing defeasible reasoning. Moreover, we want to extend the comparison to other kinds of non-monotonic logics like preferential approaches [21, 22] and modal approaches [16], but also to applications of non-monotonic logics in the context of description logics like in [8, 23]. Another interesting discussion we would like to develop is the possibility of representing defeasible reasoning through monotonic logics. Finally, we plan to exploit the analysis developed here regarding the notions of exception and generalisation to refine our formal system outlined in [10].